(Yoggo Soft Drink) A small, privately owned Asian company is producing a private-label soft drink, Yoggo. A machine-paced line puts the soft drinks into plastic bottles and then packages the bottles into boxes holding 10 bottles each. The machine-paced line is com-prised of the following four steps: (1) the bottling machine takes 1 second to fill a bottle, (2) the lid machine takes 3 seconds to cover the bottle with a lid, (3) a labeling machine takes 5 seconds to apply a label to a bottle, and (4) the packaging machine takes 4 seconds to place a bottle into a box. When a box has been filled with 10 bottles, a worker tending the packaging machine removes the filled box and replaces it with an empty box. Assume that the time for the worker to remove a filled box and replace it with an empty box is neg-ligible and hence does not affect the capacity of the line. At step 3 there are two labeling machines that each process alternating bottles; that is, the first machine processes bottles 1, 3, 5, . . . and the second machine processes bottles 2, 4, 6, . . . . Problem data are sum-marized in the table following.


a. What is the process capacity (bottles/hour) for the machine-paced line?

b. What is the bottleneck in the process?

c. If one more identical labeling machine is added to the process, how much is the increase in the process capacity going to be (in terms of bottles/hour)?

d. What is the implied utilization of the packaging machine if the demand rate is 60 boxes/ hour? Recall that a box consists of 10 bottles.
Answer #1

Output of station 1 = 60 /1 = 60 per minutes

Output of station 2 = 60/3 =20 per minute

Output of station 3 = 2x 60/5 =24 per minute

Output of station 4 = 60/4 =15 per min

(i) Process capacity of the machine paced line is 15 per min or 15x60=900 bottles per hour

(ii) Station 4 is bottleneck operation.

(iii) If one labelling machine is added, there will be no effect on the process capacity bcause labelling is not the bottleneck.

(iv) Implied utilisation of packaging = utilised capacity / maximum capacity = 60 x 10 /900 =0.6666

=66.66%